# Transformer-based Wireless Symbol Detection Over Fading Channels

### v1

@@ Abstract
Pre-trained Transformers, through in-context learning (ICL), have demonstrated exceptional capabilities to adapt to new tasks using example prompts without model update. Transformer-based wireless receivers, where prompts consist of the pilot data in the form of transmitted and received signal pairs, have shown high detection accuracy when pilot data are abundant. However, pilot information is often costly and limited in practice. In this work, we propose the DEcision Feedback IN-ContExt Detection (DEFINED) solution as a new wireless receiver design, which bypasses channel estimation and directly performs symbol detection using the (sometimes extremely) limited pilot data. The key innovation in DEFINED is the proposed decision feedback mechanism in ICL, where we sequentially incorporate the detected symbols into the prompts as pseudo-labels to improve the detection for subsequent symbols. Furthermore, we proposed another detection method where we combine ICL with Semi-Supervised Learning (SSL) to extract information from both labeled and unlabeled data during inference, thus avoiding the errors propagated during the decision feedback process of the original DEFINED. Extensive experiments across a broad range of wireless communication settings demonstrate that a small Transformer trained with DEFINED or IC-SSL achieves significant performance improvements over conventional methods, in some cases only needing a single pilot pair to achieve similar performance of the latter with more than 4 pilot pairs.
@@
Symbol Detection; Decision Feedback; MIMO; In-Context Learning; Transformer; Semi-Supervised Learning.
@@ I Introduction
Wireless receiver symbol detection focuses on identifying the transmitted symbols over fading channels with varying signal-to-noise ratios (SNRs). Traditional methods typically follow a two-step process: first estimating the channel using, e.g., the Minimum Mean Square Error (MMSE) estimator, and then performing symbol detection using the estimated channel. However, this approach can be computationally intensive when the system dimension is high, and is highly dependent on the quality of channel estimation. With the advance of machine learning (ML), data-driven approaches, such as deep learning models that directly learn channel estimators and symbol detectors, offer an alternative. Compared to the prior model-based, analytically driven solutions, these ML approaches have better adaptability to complex environments that have model mismatches, handle nonlinearity and system imperfections more effectively, and allow for joint optimization of multiple modules (or even the entire transceiver chain).
Despite their potential, deep learning (DL) based solutions have not been widely deployed in real-world wireless systems. Among the many challenges, data dependency and generalization issues stand out as one of the most significant obstacles [1]. DL models require large amounts of diverse and high-quality training data to train, which is difficult to obtain in many wireless deployments. Poor generalization to new, unseen wireless environments (e.g., new channel conditions, interference patterns) is a particularly significant challenge due to the diverse use cases in wireless systems. Although meta-learning [2] can help with adapting to new tasks, existing approaches [3, 4] often rely on explicit model parameter updates, which increase computational complexity and reduce robustness during adaptation.
Advances in Transformer models [5], particularly decoder-only architectures like GPT [6], have driven remarkable progress in natural language processing and a broad spectrum of other domains. Recent work [7] demonstrates that pre-trained Transformers can generalize to new tasks during inference through in-context learning (ICL), without requiring explicit model updates. Specifically, the input is structured as $(y_{1},f(y_{1}),\ldots,y_{n},f(y_{n}),y_{\text{query}})$, where $(y_{1},\ldots,y_{n})$ represent features in the problem domain and $f$ is an unknown mapping. By leveraging the Transformer’s advanced next-token prediction capability, a well-pretrained model can approximate $f(y_{\text{query}})$ with high accuracy for many classes of functions, conditioned on the given context. This capability has been further validated in [8], highlighting the flexibility of Transformers in handling unseen tasks during inference.
@@
Inspired by decision feedback in wireless communication (e.g., decision feedback equalizers over multi-path fading channels), we propose a novel prompt design by incorporating decision pairs. Specifically, we combine the current received signal with the model’s detected symbol in a pair, merging them with previous prompts to form a new, larger prompt for subsequent symbol detection. Our DEFINED model employs a carefully designed mixture training process to achieve high performance with limited pilots (sometimes as few as a single pilot) while maintaining accuracy when sufficient pilots are available. Additionally, we frame the problem from a semi-supervised learning (SSL) perspective by developing a new IC-SSL model. This model demonstrates that Transformers can extract information from both labeled pairs and unlabeled features, enabling in-context semi-supervised learning during inference with both supervised data pairs and partially observed data in the prompt. This result highlights the promising potential of Transformers in future wireless communication systems. Extensive experiments across various modulation schemes and channel distributions validate the effectiveness of our approach. We summarize our contributions as follows:
  • We propose a novel Transformer model that jointly performs channel estimation and symbol detection. Our key innovation in the DEFINED model is the incorporation of decision feedback to allow pseudo-labels to be used in conjunction with the limited pilot data, to gradually increase the effective context length and thus improve the output performance.
@@
  • We propose an IC-SSL detector as an alternative of DEFINED. IC-SSL demonstrates that Transformers can extract information from both labeled pairs and unlabeled features to make accurate detections. This highlights the ability of Transformers to perform ICL from partial observations, which showcases its great potential for wireless applications.
  • We conduct extensive experiments across various modulation schemes in both SISO and MIMO systems under varying configurations, where we show that a very small Transformer model trained by DEFINED and IC-SSL can achieve strong accuracy and generalization performance.
The remainder of this paper is organized as follows. Section II reviews the related work. Section III presents the system model and canonical methods. Section IV introduces in-context learning-based symbol detection. The proposed DEFINED method, including model structure and training details, is described in Section V, with an IC-SSL enhancement detailed in Section VI. Section VII presents experimental results, and Section VIII concludes the paper.
@@ II Related Work
Machine learning has become a crucial tool for wireless system designs [13, 14, 15, 16, 17, 18]. For channel estimation and symbol detection, [19, 20] propose variational autoencoders (VAEs), while deep learning architectures such as Convolutional Neural Network (CNN) and Recurrent Neural Network (RNN) have been explored to improve channel estimation [21, 22, 23]. CNN-based models effectively approximate the MMSE estimators [22], and end-to-end CNN frameworks jointly perform channel estimation and symbol detection [24].
The Transformer model [5] has become fundamental across various fields. GPT-2 extends Transformers for multitask learning [6], while GPT-3 introduces in-context learning (ICL), allowing adaptation solely through contextual inputs [7]. Unlike meta-learning, ICL does not require parameter updates, making it well-suited for wireless applications. Theoretical studies reveal that Transformers approximate regression tasks [8] and perform Bayesian inference [25]. Pre-training on diverse tasks enhances ICL performance [26], and Transformers exhibit implicit gradient descent during ICL, optimizing predictions without updating parameters [27, 28]. Research also highlights the importance of input structure, showing that distribution, label space, and structural cues significantly influence ICL performance [29, 30].
ICL has been applied to wireless tasks such as channel estimation and resource allocation. [9] first introduced Transformers for symbol detection, demonstrating their effectiveness as in-context estimators. [10] extended this to MIMO systems, highlighting threshold behavior in ICL with increasing pre-training tasks, while [11] explored multi-user cell-free MIMO systems. Large language models (LLMs) have also been adapted for wireless applications, including resource management and network configuration. [12] applies ICL-based LLMs for symbol detection, while [31, 32] propose WirelessLLM and WirelessAgent frameworks, respectively, leveraging knowledge alignment and multimodal data fusion. Unlike these approaches, our work focuses on training a Transformer-based receiver from scratch. Additionally, Transformers have also been used for radio fingerprinting [33], automatic modulation recognition (AMR) [34], spectrum sensing [35], and neuromorphic designs aimed at reducing power consumption [36]. Contrastive semi-supervised learning has improved AMR with fewer labeled samples [37].
@@ III-A Wireless Model
We focus on the block-fading channel model [38] in this paper, illustrated in Figure 1. In this model, the channel $H_{t}$ remains constant over a coherent time period of $T$ time slots, and is i.i.d. across different coherence periods. In other words, $H_{t}=h_{l},\forall t=(l-1)T+1,\cdots,lT$, for the $l$-th coherence period where $h_{l}$ is drawn i.i.d. from $P_{H}$. Correspondingly, the data transmission is organized into frames, where each frame has a length that is at most $T$. The frame structure is designed such that the first $k$ transmitted symbols are known and pre-determined pilot symbols, whose original purposes include assisting the receiver to perform channel estimation [39] of $h_{l}$ so that it can perform coherent symbol detection. Generally speaking, based on the reception of a few pilot pairs $D_{k}=\{(y_{1},x_{1}),\cdots,(y_{k},x_{k})\}$, The goal is to design a demodulator that accurately recovers the transmitted symbols $x_{k+1},\cdots,x_{T}$ from the received signals $y_{k+1},\cdots,y_{T}$ with high probability while maintaining reasonable complexity.
@@ IV In-Context Learning-Based Symbol Detection
In this section, we introduce the ICL-based symbol detection method. We first formulate the symbol detection problem, and then present the ICL implementation using a popular Transformer model GPT-2, which also serves as the backbone of our proposed solution [^1] We use GPT-2 with elaborated design choices as a concrete example throughout the paper. However, the proposed principle can be easily adapted to other Transformer architectures..
IV-A ICL-based Problem Formulation
Each ICL symbol detection task $\tau$ corresponds to a latent channel $H$ and a channel noise level $\sigma^{2}$, drawn from the unknown joint distribution $P_{\tau}=P_{H}P_{\sigma^{2}}$. The ICL-based symbol detection does not have prior knowledge of the specific task $\tau$, meaning it does not know the current channel realization $H$ or the SNR level $1/\sigma^{2}$. Instead, it is only provided only with a prompt $S_{t}^{\tau}=(D_{k}^{\tau},y_{t})$, consisting of $k$ target pairs $D_{k}^{\tau}=\{(y_{1},x_{1}),\cdots,(y_{k},x_{k})\}$, which are sampled from the conditional distribution $P_{x,y|\tau}$ and serve as the in-context examples for the current task $\tau$, along with the query $y_{t}$, $\forall t=k+1,\cdots,T$.
As previously mentioned, each context pair $(x_{i},y_{i})$ and the query pair $(y_{t},x_{t})$ are i.i.d. samples given task $\tau$. For block-fading channels, the context set $D_{k}^{\tau}$, also referred to as pilot signals in wireless communication, enhances the channel estimation, thereby improving the reliability and accuracy of data transmission. We note that a significant advantage of the proposed solution is that there is no need to change the existing frame structure or the design of pilot signals. Rather, the innovation is entirely at the receiver side where we leverage the pilot and decoded signals in a different way. This is an important advantage in practice as it allows for (backward) compatibility with the existing standard.
The goal of symbol detection is to identify the corresponding input signal $x_{t}$ for the new query signal $y_{t}$ from the same task. The ICL-based detection makes its decision as $\hat{x}_{t}=f_{\theta}(S_{t}^{\tau})$, where $\theta$ represents the parameters of the model. The detection for the query is measured by the Symbol Error Rate (SER), which is the frequency at which transmitted symbols are incorrectly decoded. The expected SER for the new query with $k$ contexts, taking the expectation over the task distribution for $\forall k=1,\cdots,T-1$, is defined as:
@@ IV-B Vanilla In-Context Symbol Detection
Transformer models have emerged as powerful tools for classification tasks [42], leveraging their ability to capture long-range dependencies. This approach to wireless symbol detection was introduced in [9, 10]. The input-output structure is illustrated in Figure 2. With a causal masked self-attention mechanism, the model outputs the detection $\hat{x}_{t}$ at the corresponding position of $y_{t}$, relying only on known preceding contexts and the received signal. During the forward process, the Transformer solves $k+1$ detection problems for the same task $\tau$, using an increasing number of pilot data points. The results in [9, 10] demonstrate that the Transformer exhibits strong capabilities in symbol estimation within context, without requiring explicit model updates.
@@ V Decision Feedback In-Context Detection
The vanilla ICL approaches for symbol detection require sufficient context to achieve accurate estimation, which is often impractical in real-world scenarios. Pilot signals are costly and limited, reducing their adaptability for practical applications. Furthermore, adding more pilot signals reduces the amount of slots for data transmission, effectively reducing the system throughput. For situations where the number of pilots is small, neither conventional two-stage (channel estimation then symbol detection) nor vanilla ICL solutions can achieve good performance. Additionally, these approaches generally formulate the symbol detection task as a regression problem, as in [26, 10, 11], where a Transformer model is trained to minimize the MSE loss. Although their models achieve performances comparable to the optimal MMSE estimator for $x$, an additional projection step is required to map the output to the appropriate transmitted symbol, leading to a mismatch and losing optimality in the process.
In contrast, we directly formulate the problem as a multi-class classification task [42], allowing the model to perform symbol detection while directly optimizing for SER during inference. This approach eliminates the need for explicit channel estimation, as the model implicitly gains knowledge of the channel through the pilot signal pairs. Additionally, we generalize the approach to effectively handle scenarios where pilot information is highly limited by sequentially feeding back the already decoded symbol pairs as noisy pilots and incorporating them as part of the prompt. Our model demonstrates robust performance even in challenging conditions with only a single pilot, outperforming previous ICL models that struggle with insufficient pilot data. At the same time, it maintains high accuracy when sufficient pilot data is available.
@@
Inspired by the decision-feedback concept in wireless communication, we propose the DEcision Feedback IN-ContExt Detection (DEFINED) method for symbol detection, as illustrated in Figure 3. The key innovation is that the DEFINED model extends the prompt by incorporating the previously received signals and detection decisions alongside prior prompts to improve subsequent detections. While traditional decision-feedback equalizers (DFEs) address inter-symbol interference (ISI), our study focuses on narrowband channels without ISI. Nevertheless, decision feedback is effective here due to the latent common channel. Noisy feedback also provides valuable information, further improving model detection.
@@ V-A Tokenization, Attention, and Detection
We delve into the implementation of the Transformer for symbol detection, focusing on signal processing and its most significant components, the self-attention mechanism, as illustrated in Figure 4. The computation involves several key steps.
First, tokenization converts the input IQ samples into embeddings, which are then processed through $L$ cascaded layers of the decoder backbone, adapted from the GPT-2 model. The model employs masked multi-head self-attention, a pivotal operation in Transformer architectures. This attention mechanism enables the output embeddings to capture long-range dependencies and extract relevant information from the input sequence. In the context of symbol detection, these dependencies represent the relationships between earlier pilot pairs and the current detection task involving the received query signal, both of which depend on the same channel $h_{l}$ in the block fading model. Capturing these relationships is critical for learning signal characteristics and the underlying channel structure, ultimately improving detection accuracy.

### v2

# Decision Feedback In-Context Learning for Wireless Symbol Detection

@@ Abstract
Pre-trained Transformers, through in-context learning (ICL), have demonstrated exceptional capabilities to adapt to new tasks using example prompts without model update. Transformer-based wireless receivers, where prompts consist of the pilot data in the form of transmitted and received signal pairs, have shown high detection accuracy when pilot data are abundant. However, pilot information is often costly and limited in practice. In this work, we propose DEcision Feedback IN-ContExt Detection (DEFINED) as a new wireless receiver design, which bypasses channel estimation and directly performs symbol detection using the (sometimes extremely) limited pilot data. The key innovation in DEFINED is the proposed decision feedback mechanism in ICL, where we sequentially incorporate the detected symbols into the prompts as pseudo-labels to improve the detection for subsequent symbols. We further establish an error lower bound and provide theoretical insights into the model’s generalization under channel distribution mismatch. Extensive experiments across a broad range of wireless settings demonstrate that a small Transformer trained with DEFINED achieves significant performance improvements over conventional methods, in some cases only needing a single pilot pair to achieve similar performance to the latter with more than 4 pilot pairs.
@@
In-Context Learning, Transformer, Symbol Detection, Decision Feedback, Wireless Channels
@@ I Introduction
Wireless receiver symbol detection aims to recover transmitted symbols over fading channels with varying SNRs. Traditional methods use a two-step process: channel estimation (e.g., via Minimum Mean Square Error (MMSE)) followed by symbol detection. However, this approach can be computationally intensive and sensitive to estimation errors, especially in high-dimensional systems. In contrast, machine learning-based approaches, including deep learning models, enable end-to-end learning that better adapts to complex, nonlinear environments and system imperfections, while supporting joint optimization across transceiver components. Such data-driven designs align with the vision of AI-native 6G systems [2], where learning-based architectures are integrated from the ground up to enhance adaptability, efficiency, and performance.
Despite their potential, deep learning (DL) based solutions have not been widely deployed in real-world wireless systems. Among the many challenges, data dependency and generalization are among the most significant obstacles [3]. DL models require large amounts of diverse and high-quality training data to train, which is difficult to obtain in many wireless deployments. Poor generalization to new, unseen wireless environments (e.g., new channel conditions, interference patterns) is a particularly significant challenge due to the diverse use cases in wireless systems. Although meta-learning [4] can help with adapting to new tasks, existing approaches [5, 6] often rely on explicit model parameter updates, which increase computational complexity and reduce robustness during adaptation.
Advances in Transformer models [7], particularly decoder-only architectures like GPT [8], have driven remarkable progress in natural language processing and a broad spectrum of other domains. Recent work [9] demonstrates that pre-trained Transformers can generalize to new tasks during inference through in-context learning (ICL), without requiring explicit model updates. Specifically, the input is structured as $(y_{1},f(y_{1}),\ldots,y_{n},f(y_{n}),y_{\text{query}})$, where $(y_{1},\ldots,y_{n})$ represent features in the problem domain and $f$ is an unknown mapping. By leveraging the Transformer’s advanced next-token prediction capability, a well-pretrained model can approximate $f(y_{\text{query}})$ with high accuracy for many classes of functions, conditioned on the given context.
@@
Inspired by decision feedback in wireless communication (e.g., decision feedback equalizers over multi-path fading channels), we propose a novel prompt design by incorporating decision pairs. Specifically, we combine the current received signal with the model’s detected symbol in a pair, merging them with previous prompts to form a new, larger prompt for subsequent symbol detection. Our DEFINED model employs a carefully designed mixture training process to achieve high performance with limited pilots (sometimes as few as a single pilot) while maintaining accuracy when sufficient pilots are available. This result highlights the promising potential of Transformers in future wireless communication systems. Extensive experiments across various modulation schemes and channel distributions validate the effectiveness of our approach. We summarize our contributions as follows:
  • We propose a novel Transformer model that jointly performs channel estimation and symbol detection. Our key innovation in DEFINED is the incorporation of decision feedback to allow decision feedback to be used in conjunction with the limited pilot data, to gradually increase the effective context length and thus improve the output performance.
@@
  • Theoretically, we derive an error lower bound and provide theoretical insights into the model’s generalization capability under channel distribution mismatch.
  • We conduct extensive experiments across various modulation schemes under varying configurations, where we show that a very small Transformer model trained by DEFINED can achieve strong accuracy and generalization performance.
The remainder of this paper is organized as follows. Section II reviews the related work. Section III presents the system model and canonical methods. Section IV introduces in-context learning-based symbol detection. The proposed DEFINED method, including the model structure and training details, is described in Section V. Section VI provides the theoretical analysis of DEFINED. Section 8 presents experimental results, and Section VIII concludes the paper.
@@ II Related Work
Machine learning is playing a pivotal role in shaping wireless system design and is a key enabler of AI-native 6G architectures [14, 15, 16, 17, 18, 19]. The Transformer model [7] has become fundamental across various fields. GPT-2 extends Transformers for multitask learning [8], while GPT-3 introduces in-context learning (ICL), allowing adaptation solely through contextual inputs [9]. Unlike meta-learning, ICL does not require parameter updates, making it well-suited for wireless applications. Theoretical studies reveal that Transformers approximate regression tasks [20] and perform Bayesian inference [21]. Pre-training on diverse tasks enhances ICL performance [22], and Transformers exhibit implicit gradient descent during ICL, optimizing predictions without updating parameters [23, 24]. Research also highlights the importance of input structure, showing that distribution, label space, and structural cues significantly influence ICL performance [25].
ICL has been applied to wireless tasks such as channel estimation and resource allocation. [10] first introduced Transformers for symbol detection, demonstrating their effectiveness as in-context estimators. [11] extended this to MIMO systems, highlighting threshold behavior in ICL with increasing pre-training tasks, while [12] explored multi-user cell-free MIMO systems. LLMs have also been adapted for wireless applications, including resource management and network configuration. [13] applies ICL-based LLMs for symbol detection, while [26, 27] propose WirelessLLM and WirelessAgent frameworks, respectively, leveraging knowledge alignment and multimodal data fusion. Unlike these approaches, our work focuses on training a Transformer-based receiver from scratch. Additionally, Transformers have also been used for radio fingerprinting [28], automatic modulation recognition (AMR) [29], spectrum sensing [30], and neuromorphic designs aimed at reducing power consumption [31].
@@ III-A Wireless Model
We focus on the block-fading channel model [32] as illustrated in Figure 1. In this model, the channel $H_{t}$ remains constant over a coherent time of $T$ time slots, and is i.i.d. across different coherence periods. In other words, $H_{t}=h_{l},\forall t=(l-1)T+1,\cdots,lT$, for the $l$-th coherence period where $h_{l}$ is drawn i.i.d. from $P_{H}$. Correspondingly, the data transmission is organized into frames, where each frame has a length that is at most $T$. The frame structure is designed such that the first $k$ transmitted symbols are known and pre-determined pilot symbols, whose original purposes include assisting the receiver to perform channel estimation [33] of $h_{l}$ so that it can perform coherent symbol detection. Based on the reception of pilot pairs $D_{k}=\{(y_{1},x_{1}),\cdots,(y_{k},x_{k})\}$, the goal is to design a demodulator that accurately recovers the transmitted symbols $x_{k+1},\cdots,x_{T}$ from the received signals $y_{k+1},\cdots,y_{T}$ with high probability and reasonable complexity.
@@ IV In-Context Learning-Based Symbol Detection
In this section, we discuss the ICL-based symbol detection method. We first formulate the symbol detection problem, and then present the ICL implementation using a popular Transformer model GPT-2, which also serves as the backbone of our proposed solution [^1] We use GPT-2 with elaborated design choices as a concrete example throughout the paper. However, the proposed principle can be easily adapted to other Transformer architectures..
IV-A ICL-based Problem Reformulation
Each ICL symbol detection task $\tau$ corresponds to a latent channel $H$ and a channel noise level $\sigma^{2}$, drawn from the unknown joint distribution $P_{\tau}=P_{H}P_{\sigma^{2}}$. The receiver does not have any prior knowledge of the specific task $\tau$, meaning it does not know the current channel realization $H$ or the SNR level $1/\sigma^{2}$. Instead, it is provided only with a prompt $S_{t}^{\tau}=(D_{k}^{\tau},y_{t})$, consisting of $k$ target pairs $D_{k}^{\tau}=\{(y_{1},x_{1}),\cdots,(y_{k},x_{k})\}$, which are sampled from the conditional distribution $P_{x,y|\tau}$ and serve as the in-context examples for the current task $\tau$, along with the query $y_{t}$, $\forall t=k+1,\cdots,T$. For block-fading channels, the context set $D_{k}^{\tau}$ corresponds to the pilot signals.
We note that a significant advantage of this reformulation (and the subsequently proposed solution) is that there is no need to change the existing frame structure or the design of pilot signals. Rather, the innovation is entirely at the receiver side where we leverage the pilot and decoded signals in a different way. This is an important advantage in practice as it allows for backward compatibility with the existing standard.
The goal of symbol detection is to identify the corresponding input signal $x_{t}$ for the new query signal $y_{t}$ from the same task due to the block-fading nature. The ICL-based detection makes its decision as $\hat{x}_{t}=f_{\theta}(S_{t}^{\tau})$, where $\theta$ represents the parameters of the model. The detection for the query is measured by the Symbol Error Rate (SER), which is the frequency at which transmitted symbols are incorrectly decoded. The expected SER for the new query with $k$ contexts, taking the expectation over the task distribution for $\forall k=1,\cdots,T-1$, is defined as:
@@ IV-B Vanilla In-Context Symbol Detection
Transformer models have emerged as powerful tools for classification tasks [41], leveraging their ability to capture long-range dependencies. This approach to wireless symbol detection was introduced in [10, 11]. The input-output structure is illustrated in Figure 2. With a causal masked self-attention mechanism, the model outputs the detection $\hat{x}_{t}$ at the corresponding position of $y_{t}$, relying only on known preceding contexts and the received signal. During the forward process, the Transformer solves $k+1$ detection problems for the same task $\tau$, using increasing pilot data points. The results in [10, 11] demonstrate that the Transformer exhibits strong capabilities in symbol estimation within context, without requiring explicit model updates.
@@ V Decision Feedback In-Context Detection
The vanilla ICL approaches for symbol detection require sufficient context to achieve accurate estimation, which is often impractical in real-world scenarios. Pilot signals are costly and limited, reducing their adaptability for practical applications. Furthermore, transmitting more pilot signals reduces system throughput and energy efficiency. For situations where the number of pilots is small, neither conventional two-stage (channel estimation then symbol detection) nor vanilla ICL solutions can achieve good performance. Additionally, the ICL approaches generally formulate the symbol detection task as a regression problem, as in [22, 11, 12], where a Transformer model is trained to minimize the MSE loss. Although their models achieve performances comparable to the optimal MMSE estimator for $x$, an additional projection step is required to map the output to the transmitted symbol, leading to a mismatch and losing optimality in the process.
In contrast, we directly formulate the problem as a multi-class classification task [41], allowing the model to perform symbol detection while directly optimizing for SER during inference. This approach eliminates the need for explicit channel estimation, as the model implicitly gains knowledge of the channel through the pilot signal pairs. Additionally, inspired by the decision-feedback concept in wireless communication, we propose the DEcision Feedback IN-ContExt Detection (DEFINED) method for symbol detection, as illustrated in Figure 3. In this approach, DEFINED sequentially feeds back the already decoded symbol pairs as noisy pilots and incorporates them as part of the prompt. As will be seen, DEFINED achieves robust performance even under challenging conditions with only a single pilot, outperforming previous ICL models that struggle with insufficient pilot data. At the same time, it maintains high accuracy when sufficient pilot data is available.
@@ V-A Tokenization, Attention, and Detection
The overall model architecture of DEFINED is given in Figure 4. The computation involves several key steps.
First, tokenization converts the input IQ samples into embeddings, which are then processed through $L$ cascaded layers of the decoder backbone that is adapted from the GPT-2 model. The model employs masked multi-head self-attention, a pivotal operation in Transformer architectures. This attention mechanism enables the output embeddings to capture long-range dependencies and extract relevant information from the input sequence. In the context of symbol detection, these dependencies represent the relationships between earlier pilot pairs and the current detection task involving the received query signal, both of which depend on the same channel $h_{l}$ in the block fading model. Capturing these relationships is critical for learning signal characteristics and the underlying channel structure, ultimately improving detection accuracy.